# Structured and Balanced Multi-component and Multi-layer Networks

### v1

@@ 4.2 Highly oscillatory functions
Globally oscillatory functions with significant high-frequency components can not be approximated well by a shallow network when a global bounded activation function of the form $\sigma({\bm{W}}\cdot{\bm{x}}-{\bm{b}})$, such as ReLU, is used. Due to almost orthogonality or high decorrelation (in terms of the inner product) between $\sigma({\bm{W}}\cdot{\bm{x}}-{\bm{b}})$ and oscillatory functions with high likelihood (in terms of a random choice of $({\bm{W}},{\bm{b}})$), the set of parameters that can render a good approximation, namely the Rashomon set [29], becomes smaller and smaller (in terms of relative measure) and hence harder and harder to find as the target function becomes more and more oscillatory (see [38]). Although this difficulty can be alleviated by complexity decomposition using MMNN as shown in Section 2, it still requires a larger network in terms of width, rank, and layers and more training. Here we limit our tests to oscillatory functions in 1D and 2D due to the dramatic increase of complexity with dimensions, or the curse of dimensions, in general.
We again start with a 1D example, $f(x)=\sin(50\pi x),x\in[-1,1]$. A MMNN of size $(800,40,15)$ produces a good approximation of this highly oscillatory function, as illustrated by the error plot in Figure 20, with a smaller learning rate and a longer training process compared to previous examples with localized fine features. Due to the significant depth, we consider using ResMMNN as discussed in Section 2.2. For this test, a total of $1000$ uniformly sampled points in $[-1,1]$ are used with a mini-batch size of $100$ and a learning rate of $10^{-4}\times 0.9^{\lfloor k/800\rfloor}$, where $k=1,2,\mathinner{\mkern-0.1mu\cdotp\mkern-0.3mu\cdotp\mkern-0.3mu\cdotp\mkern-0.1mu},40000$ is the epoch number. Also, an interesting learning dynamics for Adam is observed from Figure 20. In the beginning, nothing seems to happen until about epoch 3600 when learning starts from the boundary. Then more and more features are captured from the boundary to the inside gradually. Eventually, all features are captured and then fine-tuned together to improve the overall approximation.
@@
Figure 20: Illustrations of training and test errors measured in MSE.
@@
$(a_{i,j})=\begin{bmatrix}0.3&0.2\\ 0.2&0.3\end{bmatrix},\qquad(b_{i})=\begin{bmatrix}2\pi\\ 4\pi\end{bmatrix},\qquad(c_{i,j})=\begin{bmatrix}2\pi&4\pi\\ 8\pi&4\pi\end{bmatrix},\quad\textnormal{and}\quad(d_{i,j})=\begin{bmatrix}4\pi&6\pi\\ 8\pi&6\pi\end{bmatrix}.$
In our test, we choose $s=3$ to ensure the function exhibits significant oscillations and contains diverse Fourier modes as illustrated by Figure 22. Given the complexity of the function, we employ a MMNN with size $(600,30,15)$. Again, ResMMNN is used due to the depth. For this test, a total of $400^{2}$ data are sampled on a uniform grid in $[-1,1]^{2}$ with a mini-batch size of $1000$ and a learning rate of $10^{-3}\times 0.9^{\lfloor k/40\rfloor}$, where $k=1,2,\mathinner{\mkern-0.1mu\cdotp\mkern-0.3mu\cdotp\mkern-0.3mu\cdotp\mkern-0.1mu},2000$ is the epoch number. The training process is illustrated by Figure 23. Figure 22 shows $\log$-error plot.
@@
Figure 22: Training and test errors measured in MSE.
@@
Figure 23: The top row: the learned neural network; the bottom row: the differences between the learned neural network and the target function.
@@
(a) Approximation by MMNN.
(b) Difference.
(c) Errors (in MSE) vs. epoch
Figure 24: Approximation in a unit disk.
4.3 Tests in three dimension and higher
where
Figure 25: Surface plot of the levelset $f(r,\theta,\phi)=0.5$.
Figure 26: Surface plot of the levelset $h(r,\theta,\phi)=0.5$.
Figure 27: Training and test errors (MSE) vs. epoch.
Our MMNN is of a compact size $(600,20,8)$. For this test, a total of $111^{3}$ data are sampled on a uniform grid in $[-1,1]^{3}$ with a mini-batch size of $999$ and a learning rate of $0.0005\times 0.9^{\lfloor k/6\rfloor}$ for epochs $k=1,2,\mathinner{\mkern-0.1mu\cdotp\mkern-0.3mu\cdotp\mkern-0.3mu\cdotp\mkern-0.1mu},300$. Figure 27 gives the error plot. As shown in Figures 27 and 27, the levelsets corresponding to the target function $f$ and the learned MMNN approximation $h$ are nearly identical. To visually demonstrate the quality of the approximation and complex structure of the 3D function, we present several slices of the target function and the MMNN approximation by fixing either $x$, $y$, or $z$ in Figure 28.
(a) $z=0$.
(b) $z=0.1$.
(c) $y=0.2$.
(d) $y=0.3$.
(e) $x=0.4$.
(f) $x=0.5$.
$f(\bm{x})=f(x_{1},\dots,x_{4})=\frac{\exp\left(-\frac{1}{2}(\bm{x}-\bm{\mu})^{\textsf{T}}{\bm{\Sigma}}^{-1}(\bm{x}-\bm{\mu})\right)}{\sqrt{(2\pi)^{k}\det({\bm{\Sigma}})}}$
where ${\bm{\Sigma}}$ is the covariance matrix. We set $\bm{\mu}=\bm{0}$ and ${\bm{\Sigma}}^{-1}=20\left[\begin{smallmatrix}1.0&0.9&0.8&0.7\\ 0.9&2.0&1.9&1.8\\ 0.8&1.9&3.0&2.9\\ 0.7&1.8&2.9&4.0\\ \end{smallmatrix}\right].$ We remark that the eigenvalues of ${\bm{\Sigma}}^{-1}$ are $6.82,9.93,25.28,158.05$ which means that the distribution is quite anisotropic and concentrated near the center.
Figure 30: Training and test errors (MSE) vs. epoch.
@@ 4.4 Learning dynamics
We again start with a 1D example, $f(x)=\sin\big{(}36\pi|x|^{1.5}\big{)},x\in[-1,1]$. A MMNN of size $(600,30,8)$ produces a good approximation of this highly oscillatory function, as illustrated by the error plot in Figure 32. For this test, a total of $1000$ uniformly sampled points in $[-1,1]$ are used with a mini-batch size of $100$ and a learning rate of $10^{-3}\times 0.9^{\lfloor k/200\rfloor}$, where $k=1,2,\mathinner{\mkern-0.1mu\cdotp\mkern-0.3mu\cdotp\mkern-0.3mu\cdotp\mkern-0.1mu},10000$ is the epoch number. As illustrated in Figure 32, the function is less oscillatory near 0. Therefore, we might anticipate that the network will initially learn the part near 0 and then feature by feature from the middle to the boundary. The experimental results presented in Figure 33 agree with our expectations.
Figure 31: Derivative of the target function.
@@
Now we show an example of 2D function $f(r,\theta)$ (see Figure 35) defined in polar coordinates $(r,\theta)$ as
@@
Figure 34: Illustration of the target function.
Figure 35: Training and test errors (in MSE) vs. epoch.
@@ 5 Further discussion
In this section, we provide several general insights about MMNNs. First, in Section 5.1, we explore the advantages of MMNNs compared to fully connected networks (FCNNs) or multi-layer perceptrons (MLPs). Next, in Section 5.2, we offer practical guidelines for determining the appropriate MMNN size based on our theoretical understanding and extensive numerical experiments. Finally in Section 5.3, we discuss the use of alternative activation functions beyond $\mathtt{ReLU}$ in MMNNs.
@@ 5.1 Advantages compared to FCNNs or MLPs
  • The two key differences between a standard FCNN or MLP and a MMNN are 1) the introduction of the weights ${\bm{A}},{\bm{c}}$ for different linear combinations of hidden neurons (or perceptrons) as the multi-components in each layer, and 2) the training strategy that fixes those randomly initialized ${\bm{W}},{\bm{b}}$ (random features) in the hidden neurons. Hence it is extremely easy to modify a FCNN or MLP to a MMNN.
  • MMNNs are much more effective than FCNNs in terms of representation, training, and accuracy especially for complex functions. In comparison, as shown in those experiments in Section 2.4, MMNNs 1) have much fewer training parameters, 2) converge much faster in training, 3) achieve much better accuracy. Moreover, experiments show that training process of MMNNs converges not only faster but also with a steady rate while FCNNs saturates pretty early to a quite low accuracy, as commonly observed in practices. These nice behaviors of MMNNs are due to their balanced structure for smooth decomposition as well as the training strategy. In practice, the introduction of ${\bm{A}},{\bm{c}}$ in MMNNs provides an important balance between the network width, which is the number of hidden neurons (basis functions) and can be very large, and the dimension of the input space, which is the number of components from the previous layer and can be much smaller than the network width. In other words, using a few linear combinations of the basis functions can capture smooth structures in the input space well. On the other hand, for FCNNs the two are the same and no balance is exerted.
@@ 5.2 Practical guidelines for MMNN
  1. 1.
    The network width should provide enough resolution to capture fine details of the target function. This means that the width should be at least comparable to the size of an adaptive mesh that can approximate the target function well.
  2. 2.
    The number of components (rank) is related to the overall complexity of the target function which depends on its spatial domain and Fourier domain representation as well as the input dimension. As indicated by our mathematical multi-component construction, it is related to the “divide and conquer” strategy.
  3. 3.
    The number of layers (depth) is also related to the overall complexity of the target function as for the number of components. Rank and depth are complementary but work together effectively for a smooth decomposition of the target function. The rule of thumb for depth is similar to that for the rank.
@@
  • Overall, for a given target function, MMNNs can work well with quite a large range of configuration with a trade-off between the network size and training process. For example, the training process for a network more on the compact size with respect to the complexity of a given target function may become more subtle and challenging, e.g., choosing the appropriate learning rate and batch size, due to the lack of flexibility (or redundancy) of the representation. On the other hand, a network of too large size (or redundancy) with respect to the complexity of a given target function requires unnecessarily expensive training cost. An interesting and important question for future research is how to develop a posteriori strategy to automatically adjust the network size in practice.
  • The most advantageous situation for using MMNNs is when approximating a function in relative high dimension which is mostly smooth except for localized fine features, e.g., a distribution in high dimensions concentrated on a low dimensional manifold. Through training, MMNNs can provide an automatic adaptive approximation of the underlying structure which can be challenging for a traditional mesh based method.
@@ 5.3 Beyond ReLU to other activation functions
We also tried using different activation functions for MMNNs, e.g., $\mathtt{GELU}$ [9], $\mathtt{Swish}$ [25], $\mathtt{Sigmoid}$, and $\mathtt{Tanh}$. In general, ReLU provides the overall best results for various target functions. However, in situations where a smooth (e.g., $C^{s}$ or real analytic) approximation is needed, one might consider using smooth alternatives to $\mathtt{ReLU}$ such as $\mathtt{GELU}$  or $\mathtt{Swish}$, which generally yield results comparable to $\mathtt{ReLU}$:
  • For target functions that are $C^{s}$ or even real analytic (and can be highly oscillatory), such as $f(x)=\cos(36\pi x^{2})-0.6\cos(12\pi x^{2})$, $\mathtt{GELU}$ (or $\mathtt{Swish}$) tends to perform slightly better than $\mathtt{ReLU}$.
  • For target functions with non-differentiable points, such as $f(x)={\mathds{1}}_{\{|x|<0.02\}}\cdot\sin(50\pi x)$, $\mathtt{GELU}$ (or $\mathtt{Swish}$) generally performs slightly worse than $\mathtt{ReLU}$.
  • The use of $\mathtt{GELU}$ (or $\mathtt{Swish}$) typically results in slightly longer training time compared to $\mathtt{ReLU}$.
@@ Acknowledgments
H. Zhao was partially supported by NSF grants DMS-2309551, and DMS-2012860. Y. Zhong was partially supported by NSF grant DMS-2309530, H. Zhou was partially supported by NSF grant DMS-2307465.

### v2

@@ 5.2 Highly oscillatory functions
Globally oscillatory functions with significant high-frequency components can not be approximated well by a shallow network when a global bounded activation function of the form $\sigma({\bm{W}}\cdot{\bm{x}}-{\bm{b}})$, such as ReLU, is used. Due to almost orthogonality or high decorrelation (in terms of the inner product) between $\sigma({\bm{W}}\cdot{\bm{x}}-{\bm{b}})$ and oscillatory functions with high likelihood (in terms of a random choice of $({\bm{W}},{\bm{b}})$), the set of parameters that can render a good approximation, namely the Rashomon set [32], becomes smaller and smaller (in terms of relative measure) and hence harder and harder to find as the target function becomes more and more oscillatory (see [41]). Although this difficulty can be alleviated by complexity decomposition using MMNN as shown in Section 2, it still requires a larger network in terms of width, rank, and layers and more training. Here we limit our tests to oscillatory functions in one-dimensional and two-dimensional cases due to the dramatic increase of complexity with dimensions, or the curse of dimensions, in general.
We again start with a one-dimensional example, $f(x)=\sin(50\pi x),x\in[-1,1]$. An MMNN of size $(800,40,15)$ produces a good approximation of this highly oscillatory function, as illustrated by the error plot in Figure 21, with a smaller learning rate and a longer training process compared to previous examples with localized fine features. Due to the significant depth, we consider using ResMMNN as discussed in Section 2.1. For this test, a total of $1000$ uniformly sampled points in $[-1,1]$ are used with a mini-batch size of $100$ and a learning rate of $10^{-4}\times 0.9^{\lfloor k/800\rfloor}$, where $k=1,2,\mathinner{\mkern-0.1mu\cdotp\mkern-0.3mu\cdotp\mkern-0.3mu\cdotp\mkern-0.1mu},40000$ is the epoch number. Also, an interesting learning dynamics for Adam is observed from Figure 21. In the beginning, nothing seems to happen until about epoch 3600 when learning starts from the boundary. Then more and more features are captured from the boundary to the inside gradually. Eventually, all features are captured and then fine-tuned together to improve the overall approximation.
@@
Figure 21: Training and test errors in MSE vs. epoch.
@@
$(a_{i,j})=\begin{bmatrix}0.3&0.2\\ 0.2&0.3\end{bmatrix},\qquad(b_{i})=\begin{bmatrix}2\pi\\ 4\pi\end{bmatrix},\qquad(c_{i,j})=\begin{bmatrix}2\pi&4\pi\\ 8\pi&4\pi\end{bmatrix},\quad(d_{i,j})=\begin{bmatrix}4\pi&6\pi\\ 8\pi&6\pi\end{bmatrix}.$
In our test, we choose $s=3$ to ensure the function exhibits significant oscillations and contains diverse Fourier modes as illustrated by Figure 23. Given the complexity of the function, we employ an MMNN with size $(600,30,15)$. Again, ResMMNN is used due to the depth. For this test, a total of $400^{2}$ data are sampled on a uniform grid in $[-1,1]^{2}$ with a mini-batch size of $1000$ and a learning rate of $10^{-3}\times 0.9^{\lfloor k/40\rfloor}$, where $k=1,2,\mathinner{\mkern-0.1mu\cdotp\mkern-0.3mu\cdotp\mkern-0.3mu\cdotp\mkern-0.1mu},2000$ is the epoch number. The training process is illustrated by Figure 24. Figure 23 shows $\log$-error plot.
@@
Figure 23: Training and test errors in MSE vs. epoch.
@@
Figure 24: The top row: the learned neural network; the bottom row: error.
@@
(a) Approximation.
(b) Error.
(c) Errors (in MSE).
Figure 25: MMNN approximation in a unit disk.
5.3 Discontinuous functions with porous structures
Earlier, we tested highly oscillatory functions that are still continuous. To be more inclusive in our experiments, we now consider piecewise smooth functions with porous structures and complicated interfaces. Two target functions are shown in Figure 26(a,c). The first is a constant function with holes of various shapes removed (piecewise constant), while the second is based on the function in Figure 4 with holes introduced (piecewise smooth). We choose an MMNN of size (256, 12, 6), denoted MMNN1, and another of size (1024, 32, 6), denoted MMNN2, to learn these functions, respectively. For training, we sample $600^{2}$ data points on a uniform grid in $[-1,1]^{2}$, using a mini-batch size of 1000 and a learning rate of $0.001\times 0.9^{\lfloor k/20\rfloor}$ for epochs $k=1,2,\mathinner{\mkern-0.1mu\cdotp\mkern-0.3mu\cdotp\mkern-0.3mu\cdotp\mkern-0.1mu},1600$.
(a) Truth.
(b) MMNN1.
(c) Truth.
(d) MMNN2.
Figure 26: True functions and learned networks.
As illustrated in Figure 26, MMNNs demonstrate impressive ability in localizing and capturing discontinuities, geometries, and oscillations simultaneously. In other words, MMNNs can be adaptive in both space and frequency domains!
@@ 5.4 Learning dynamics
We again start with a one-dimensional example, $f(x)=\sin\big{(}36\pi|x|^{1.5}\big{)},x\in[-1,1]$. An MMNN of size $(600,30,8)$ produces a good approximation of this highly oscillatory function, as illustrated by the error plot in Figure 28. For this test, a total of $1000$ uniformly sampled points in $[-1,1]$ are used with a mini-batch size of $100$ and a learning rate of $10^{-3}\times 0.9^{\lfloor k/200\rfloor}$, where $k=1,2,\mathinner{\mkern-0.1mu\cdotp\mkern-0.3mu\cdotp\mkern-0.3mu\cdotp\mkern-0.1mu},10000$ is the epoch number. As illustrated in Figure 28, the function is less oscillatory near 0. Therefore, we might anticipate that the network will initially learn the part near 0 and then feature by feature from the middle to the boundary. The experimental results presented in Figure 29 agree with our expectations.
Figure 27: Derivative of $f$.
@@
Now we show an example of two-dimensional function $f(r,\theta)$ (see Figure 31) defined in polar coordinates $(r,\theta)$ as
@@
Figure 30: Illustration of the target function.
Figure 31: Errors (in MSE) vs. epoch.
@@ 6 Further discussion
In this section, we provide a few more comments about MMNNs. First, in Section 6.1, we discuss the advantages of MMNNs over fully connected networks (FCNNs) or multi-layer perceptrons (MLPs). Next, in Section 6.2, we offer practical guidelines for determining the appropriate MMNN size based on our theoretical understanding and extensive numerical experiments. Finally in Section 6.3, we discuss the use of alternative activation functions beyond $\mathtt{ReLU}$ in MMNNs.
@@ 6.1 Advantages compared to FCNNs or MLPs
The two key differences between a standard FCNN or MLP and an MMNN are 1) the introduction of the weights ${\bm{A}},{\bm{c}}$ for different linear combinations of hidden neurons (or perceptrons) as the multi-components in each layer, and 2) the training strategy that fixes those randomly initialized ${\bm{W}},{\bm{b}}$ (random features) in the hidden neurons. Hence it is extremely easy to modify a FCNN or MLP to an MMNN.
MMNNs are much more effective than FCNNs in terms of representation, training, and accuracy especially for complex functions. In comparison, as shown in those experiments in Section 3, MMNNs 1) have much fewer training parameters, 2) converge much faster in training, 3) achieve much better accuracy. Moreover, experiments show that training process of MMNNs converges not only faster but also with a steady rate while FCNNs saturates pretty early to a quite low accuracy, as commonly observed in practices. These nice behaviors of MMNNs are due to their balanced structure for smooth decomposition as well as the training strategy. In practice, the introduction of ${\bm{A}},{\bm{c}}$ in MMNNs provides an important balance between the network width, which is the number of hidden neurons (basis functions) and can be very large, and the dimension of the input space, which is the number of components from the previous layer and can be much smaller than the network width. In other words, using a few linear combinations of the basis functions can capture smooth structures in the input space well. On the other hand, for FCNNs the two are the same and no balance is exerted.
@@ 6.2 Practical guidelines for MMNN
1. The network width should provide enough resolution to capture fine details of the target function. This means that the width should be at least comparable to the size of an adaptive mesh that can approximate the target function well.
2. The number of components (rank) is related to the overall complexity of the target function which depends on its spatial domain and Fourier domain representation as well as the input dimension. As indicated by our mathematical multi-component construction, it is related to the “divide and conquer” strategy.
3. The number of layers (depth) is also related to the overall complexity of the target function as for the number of components. Rank and depth are complementary but work together effectively for a smooth decomposition of the target function. The rule of thumb for depth is similar to that for the rank.
@@
Overall, for a given target function, MMNNs can work well with quite a large range of configuration with a trade-off between the network size and training process. For example, the training process for a network more on the compact size with respect to the complexity of a given target function may become more subtle and challenging, e.g., choosing the appropriate learning rate and batch size, due to the lack of flexibility (or redundancy) of the representation. On the other hand, a network of too large size (or redundancy) with respect to the complexity of a given target function requires unnecessarily expensive training cost. There is a trade-off between representation and optimization one needs to balance in practice. An important question for future research is how to develop a self-adaptive strategy to adjust the network size.
The most advantageous situation for using MMNNs is when approximating a function in relative high dimension which is mostly smooth except for localized fine features, e.g., a distribution in high dimensions concentrated on a low dimensional manifold. Through training, MMNNs can provide an automatic adaptive approximation of the underlying structure which can be challenging for a mesh based method.
@@ 6.3 Beyond ReLU to other activation functions
We also tried using different activation functions for MMNNs, e.g., $\mathtt{GELU}$ [10], $\mathtt{Swish}$ [28], $\mathtt{Sigmoid}$, and $\mathtt{Tanh}$. In general, ReLU provides the overall best results for various target functions. However, in situations where a smooth (e.g., $C^{s}$ or real analytic) approximation is needed, one might consider using smooth alternatives to $\mathtt{ReLU}$ such as $\mathtt{GELU}$  or $\mathtt{Swish}$, which generally yield results comparable to $\mathtt{ReLU}$.
@@ Acknowledgments
S. Zhang was partially supported by the start-up fund P0053092 from The Hong Kong Polytechnic University. H. Zhao was partially supported by NSF grants DMS-2309551, and DMS-2012860. Y. Zhong was partially supported by NSF grant DMS-2309530, H. Zhou was partially supported by NSF grant DMS-2307465.
Appendix A Tests in three dimension and higher
where
Figure 33: Surface of the levelset $f(r,\theta,\phi)=0.5$.
Figure 34: Surface of the levelset $h(r,\theta,\phi)=0.5$.
Figure 35: Training and test errors (MSE) vs. epoch.
Our MMNN is of a compact size $(600,20,8)$. For this test, a total of $111^{3}$ data are sampled on a uniform grid in $[-1,1]^{3}$ with a mini-batch size of $999$ and a learning rate of $0.0005\times 0.9^{\lfloor k/6\rfloor}$ for epochs $k=1,2,\mathinner{\mkern-0.1mu\cdotp\mkern-0.3mu\cdotp\mkern-0.3mu\cdotp\mkern-0.1mu},300$. Figure 35 gives the error plot. As shown in Figures 35 and 35, the levelsets corresponding to the target function $f$ and the learned MMNN approximation $h$ are nearly identical. To visually demonstrate the quality of the approximation and complex structure of the 3D function, we present several slices of the target function and the MMNN approximation by fixing either $x$, $y$, or $z$ in Figure 36.
(a) $z=0$.
(b) $z=0.1$.
(c) $y=0.2$.
(d) $y=0.3$.
(e) $x=0.4$.
(f) $x=0.5$.
$f(\bm{x})=f(x_{1},\mathinner{\mkern-0.1mu\cdotp\mkern-0.3mu\cdotp\mkern-0.3mu\cdotp\mkern-0.1mu},x_{4})=\frac{\exp\left(-\frac{1}{2}(\bm{x}-\bm{\mu})^{\textsf{T}}{\bm{\Sigma}}^{-1}(\bm{x}-\bm{\mu})\right)}{\sqrt{(2\pi)^{k}\det({\bm{\Sigma}})}}$
where ${\bm{\Sigma}}$ is the covariance matrix. We set $\bm{\mu}=\bm{0}$ and
${\bm{\Sigma}}^{-1}=20\begin{bmatrix}1.0&0.9&0.8&0.7\\ 0.9&2.0&1.9&1.8\\ 0.8&1.9&3.0&2.9\\ 0.7&1.8&2.9&4.0\\ \end{bmatrix}.$
We remark that the eigenvalues of ${\bm{\Sigma}}^{-1}$ are $6.82,9.93,25.28,158.05$ which means that the distribution is quite anisotropic and concentrated near the center.
Figure 38: Training and test errors (MSE) vs. epoch.